# Debiasing Machine Learning Predictions for Causal Inference Without Additional Ground Truth Data: “One Map, Many Trials” in Satellite-Driven Poverty Analysis

### v3

@@ Appendix I. Deriving Implications of Tweedie’s Formula for Causal Debiasing
Remark. In practice, plug-in consistency results may be compromised by the mechanics of ML model training, where training and inference set prediction distributions often, in finite datasets, show markedly different error dynamics (with training error collapsing to 0 while inference error possibly increases). This fact can induce forms of distribution shift in $\hat{Y}_{i}$ even if data at training and inference time are from the same distribution. That said, as $n\to\infty$, the concern of overfitting would be reduced with many model classes (devroye2013probabilistic).
@@
or the Berkson-style model as a debiasing perspective (carroll2006measurement; heid2004two), used in the main body:
@@
An empirical evaluation of this relationship on our poverty prediction dataset is presented in Figure A.I.1.
@@
Appendix II. Neural Model Details
The backbone of each model was a ResNet‑50, pre‐trained via unsupervised contrastive learning (MOCO) on the SSL4EO-L dataset, as distributed in the TorchGeo library (stewart2023ssl4eoldatasetsfoundationmodels). The images were rescaled following the same procedure as in the SSL4EO-L paper, with added data augmentation in the form of random horizontal and vertical flips. A final dense layer was appended to the ResNet to enable the regression task.
Training proceeded for up to 33 epochs. During the first three epochs, we froze all pretrained weights and trained only the newly added final layer. Thereafter, we unfroze the entire network and performed end‑to‑end fine‑tuning with early stopping on the validation set (patience = 5 epochs). Optimization used AdamW with weight decay $\left(\ell_{2}\right$. regularization coefficient $\left.=1\times 10^{-4}\right)$, batch size 128 and a cosine learning‐rate schedule with warm restarts; analyses use PyTorch 2.6. When training with the Ratledge objective, we set $\lambda_{r}=0$ to avoid double-counting $\ell_{2}$ regularization (weight decay already imposes an $\ell_{2}$ penalty). If using Adam (no decoupled weight decay), set $\lambda_{r}>0$ and weight decay to zero.
The only hyperparameter tuning that was performed was to find the value $\lambda_{b}$ used by the Ratledge loss. This was achieved using a grid search, the results of which are presented in Figure A.II.1.
The variables $\hat{k}$ and $\hat{m}$ for the LCC correction were calculated on the fold’s calibration set, while the values for Tweedie’s correction, $\sigma$, and the score function were estimated with the training data. For the score function, a Gaussian KDE was fitted using Scott’s rule for determining the bandwidth.
@@ Appendix II. Neural Model Details
Figure A.II.1: Results of the hyperparameter tuning for finding the best suited $\lambda_{b}$ for the Ratledge loss. Based of these results, we opted to use $\lambda_{b}=15$ as it had the least shrinkage bias $(b\approx 1)$.
Appendix III. Additional Empirical Results
Figure A.III.1: Estimated vs. true ATE across simulated datasets under each method (varying $\tau$); a 45-degree line indicates perfect calibration.
Figure A.III.2: Treatment sites.
Figure A.III.3: Performance comparing MAE and calibration slope of Tweedie and Ratledge adjustment, seen via scatterplot of observed versus predicted mean IWI for each survey in the test set. The standard MSE-trained model produces biased estimates, overestimating wealth in poorer countries and underestimating it in richer ones. Applying Tweedie’s or LCC adjustments significantly mitigates this shrinkage bias. Training a model with Ratledge loss also mitigates shrinkage bias in this context, but results in higher variance than with Tweedie adjustment (seen by lower MAE for Tweedie).
Figure A.III.4: Scatterplot of the observed versus predicted mean IWI for each Admin 2 region in the test set, similar to A.III.3. The standard MSE-trained model produces biased estimates, overestimating wealth in poorer countries and underestimating it in richer ones. Applying Tweedie’s or LCC adjustments significantly mitigates this shrinkage bias. Training a model with Ratledge loss also mitigates the shrinkage bias, but has a relatively high error in terms of MAE.
Figure A.III.5: Same as for A.III.4, but only including Admin 2 regions with more than 100 surveyed clusters. The Tweedie adjusted model is now practically unbiased (in regards to shrinkage), suggesting that the correction has larger issues with smaller samples.
Figure A.III.6: Comparison of predicted treatment effects for aid interventions funded by China versus the World Bank, highlighting differences in estimated impacts across sectors after applying debiasing corrections.
Figure A.III.7: Scatter plots comparing ATE estimated using predicted IWI versus observed IWI for aid interventions funded by China (red points) and the World Bank (blue points). Each panel evaluates a different debiasing approach applied to predictions from an MSE-trained model: (top-left) naive (uncorrected) predictions, (top-right) Tweedie’s correction, (bottom-left) LCC, and (bottom-right) Ratledge model. Dashed black lines represent fitted linear regressions, with slopes and MAE provided; a slope close to 1 indicates reduced attenuation bias toward zero.
Figure A.III.8: Performance on a downstream causal inference task as a function of upstream model training epochs. Early training (e.g., after one epoch) yields poor model accuracy (high MAE) but minimal overfitting, allowing Tweedie’s correction to produce unbiased estimates using variance from either the calibration set ($\sigma_{\text{Cal.}}$) or training set ($\sigma_{\text{Train}}$). As epochs increase, model accuracy improves, but growing overfitting causes over-correction when using $\sigma_{\text{Cal.}}$, increasing bias, while $\sigma_{\text{Train}}$ maintains near-unbiased results.
@@ Appendix III. Additional Empirical Results
The DHS displaces survey coordinates ($\geq$ 3 km) to protect privacy, so we already address a built‑in corruption source, which shows improvement. Generally, a weak method or noisy data would make debiasing more critical, as uncertainty increases shrinkage. To show this, we ran an additional sensitivity analysis on the causal aid task (cf. Fig 6). We trained our model on the same dataset but shuffled a share of the labels in the train+cal to add noise, only increasing Tweedie’s relative advantage. Results are below in Tables A.III.1 and A.III.2.
Table A.III.1: Performance of the naive estimator under varying levels of label shuffling in training and calibration data.
@@
Table A.III.2: Performance of Tweedie’s estimator under varying levels of label shuffling in training and calibration data.
@@
Variance source for Tweedie. Figure A.III.8 shows that estimating $\sigma$ on the training split (rather than the calibration split) mitigates over-correction at later epochs; our package therefore exposes both options, defaulting to $\sigma_{\text{Train}}$. We also present in A.III.3 the results of an analysis probing the sensitivity of correction performance to cal set size (mean $\pm$ sd across 20 random subsamples).
Table A.III.3: Sensitivity of LCC and Tweedie’s estimators to calibration set size (mean $\pm$ sd across 20 random subsamples).
@@
Table A.III.4: Characterization of Debiasing Methods for Causal Inference Using ML Predictions
@@
(i) PPI (angelopoulos2023prediction) Fresh ground-truth labeled data at the downstream causal inference stage (new treatment-predictor-outcome tuples). None; treats model as black-box and applies rectifier term. Predictive model is fixed (not retrained); valid statistical coverage in large samples; independence assumptions for inference.
(ii) ratledge_using_2022 None additional; uses existing upstream training data. Yes; modifies loss function to include quintile-specific bias penalty ($\mathcal{L}=\mathrm{MSE}+\lambda_{r}\|\theta\|_{2}^{2}+\lambda_{b}\hat{E}_{b}$; if using AdamW, set $\lambda_{r}=0$). Requires retraining. Bias reducible by penalizing group-level discrepancies; assumes quintile-based error structure.
(iii) DSL (dsl_original) Gold-standard labeled data on a subset of units (no new data collection); known or estimated sampling probabilities depending on surrogates and covariates. None; post-hoc construction of bias-corrected pseudo-outcomes via cross-fitting and doubly-robust estimation. Known sampling design with probabilities bounded away from zero; double robustness (consistent if either outcome model or sampling model is correctly specified).
@@
(v) Tweedie’s correction (this work) Held-out data from upstream for estimating $\sigma^{2}$ and score function (no new data collection). None; post-hoc adjustment using empirical Bayes and density scores. Gaussian homoskedastic noise ($\hat{Y}_{i}=Y_{i}-\varepsilon_{i}$, $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ independent of $Y_{i}$); $\mathbb{E}[\varepsilon_{i}\mid A_{i}]=0$.
Funder Sector Code AD Sector Name Observed ATE (IWI) Naive Predicted ATE (IWI) Tweedie’s ATE (IWI) Num. Treated Clusters Num. Control Clusters
@@
Table A.III.5: Overview: Aid sector effect estimates.
Table A.III.6: Overview of data: DHS surveys on health and living conditions.
@@
Gabon 719 43.10 16.50 12, 19-21 59.0
Gambia 280 51.17 14.75 19-20 61.8
Ghana 2456 39.11 15.77 93, 98, 03, 08, 14, 16, 19, 22 45.8
Guinea 1442 30.04 15.60 99, 05, 12, 18, 21 35.9
Kenya 4605 30.41 14.85 03, 08-09, 14, 15, 20, 22 38.9
Lesotho 1162 26.72 12.23 04, 09, 14 27.1
Liberia 1534 21.45 12.64 07, 09, 11, 13, 16, 19-20, 22 40.7
Madagascar 2393 21.25 12.86 97, 08-09, 11, 13, 16, 21 25.8
Malawi 3184 19.68 9.31 00, 04, 10, 12, 14, 15-16, 17 20.5
Mali 2182 26.90 13.48 95-96, 01, 06, 12-13, 15, 18, 21 32.5
Mauritania 1198 38.18 18.67 19-21 47.2
Morocco 476 54.56 17.11 03-04 56.7
Mozambique 1406 26.41 16.91 09, 11, 15, 18 42.7
Namibia 1301 37.29 21.03 00, 06-07, 13 44.7
Niger 899 21.52 15.04 92, 98, 12, 21 36.2
Nigeria 4641 36.78 18.08 03, 08, 10, 13, 15, 18, 21 38.7
Rwanda 2185 22.12 9.34 05, 07-08, 10, 14-15, 19-20 21.6
Senegal 3387 37.48 16.81 92-93, 97, 05, 08-09, 10-11, 12-19, 20-21 41.1
Sierra Leone 1678 24.21 13.97 08, 13, 16, 19 36.3
South Africa 746 58.23 14.43 16 62.2
Tanzania 3626 24.48 13.97 99, 03-04, 07-08, 10, 11-12, 15-16, 17, 22 27.7
Togo 621 30.83 15.18 98, 13-14, 17 37.8
Uganda 2417 24.89 13.52 00-01, 06, 09, 11, 14-15, 16, 18-19 27.2
Zambia 1573 26.73 17.96 07, 13-14, 18 39.2
Zimbabwe 1401 32.83 17.83 99, 05-06, 10-11, 15 38.1
@@
Appendix IV. Simulation Design
@@
Appendix V. World Bank and China Empirical Case
For robustness, we repeat this experiment, using urban/rural, country, and survey year as confounders (Figure A.III.6). The results show that this approach breaks the bias in terms of treatment effect. For larger studies, with more than 200 treated points, it also lowers the total error rate. We hypothesize that the reason the overall error rate does not decrease for small trials is due to the added variance introduced when calculating the Tweedie correction. The correction is a mean of the sample scores, and it decreases with larger sample sizes.
Appendix VI. Implementing Debiasing Strategies with unshrink
@@
For all-Africa IWI wealth map predictions with associated score and calibration parameters (e.g., $\sigma$) for adjustment:

### v4

@@ Introduction
Remark. In practice, plug-in consistency results may be compromised by the mechanics of ML model training, where training and inference set prediction distributions often, in finite datasets, show markedly different error dynamics (with training error collapsing to 0 while inference error possibly increases). This fact can induce forms of distribution shift in $\hat{Y}_{i}$ even if data at training and inference time are from the same distribution. That said, as $n\to\infty$, the concern of overfitting would be reduced with many model classes (Devroye, Györfi, and Lugosi 2013).
@@
or the Berkson-style model as a debiasing perspective (Carroll et al. 2006; Heid et al. 2004), used in the main body:
@@
An empirical evaluation of this relationship on our poverty prediction dataset is presented in Figure 6.
@@
Appendix B B. Neural Model Details
The backbone of each model was a ResNet‑50, pre‐trained via unsupervised contrastive learning (MOCO) on the SSL4EO-L dataset, as distributed in the TorchGeo library (Stewart et al. 2023). The images were rescaled following the same procedure as in the SSL4EO-L paper, with added data augmentation in the form of random horizontal and vertical flips. A final dense layer was appended to the ResNet to enable the regression task.
Training proceeded for up to 33 epochs. During the first three epochs, we froze all pretrained weights and trained only the newly added final layer. Thereafter, we unfroze the entire network and performed end‑to‑end fine‑tuning with early stopping on the validation set (patience = 5 epochs). Optimization used AdamW with weight decay $\left(\ell_{2}\right$. regularization coefficient $\left.=1\times 10^{-4}\right)$, batch size 128 and a cosine learning‐rate schedule with warm restarts; analyses use PyTorch 2.6. When training with the Ratledge objective with AdamW, we set $\lambda_{r}=0$ to avoid double-counting $\ell_{2}$ regularization (weight decay already imposes an $\ell_{2}$ penalty).
The only hyperparameter tuning that was performed was to find the value $\lambda_{b}$ used by the Ratledge loss. This was achieved using a grid search (Figure 15).
Quantities $\hat{k}$ and $\hat{m}$ for LCC correction were calculated on the fold’s calibration set, values for Tweedie’s correction ($\sigma$, the score function) were estimated with the training data. For scores, a Gaussian KDE was fit using Scott’s rule.
@@
Appendix C C. Additional Empirical Results
For treatment location maps by aid sector and additional figures, see: github.com/cjerzak/Figures.
Figure 7: Estimated vs. true ATE across simulated datasets under each method (varying $\tau$); a 45-degree line indicates perfect calibration.
Figure 8: Treatment sites.
Figure 9: Performance comparing MAE and calibration slope of Tweedie and Ratledge adjustment, seen via scatterplot of observed versus predicted mean IWI for each survey in the test set. The standard MSE-trained model produces biased estimates, overestimating wealth in poorer countries and underestimating it in richer ones. Applying Tweedie’s or LCC adjustments significantly mitigates this shrinkage bias. Training a model with Ratledge loss also mitigates shrinkage bias in this context, but results in higher variance than with Tweedie adjustment (seen by lower MAE for Tweedie).
Figure 10: Scatterplot of the observed versus predicted mean IWI for each Admin 2 region in the test set, similar to 9. The standard MSE-trained model produces biased estimates, overestimating wealth in poorer countries and underestimating it in richer ones. Applying Tweedie’s or LCC adjustments significantly mitigates this shrinkage bias. Training a model with Ratledge loss also mitigates the shrinkage bias, but has a relatively high error in terms of MAE.
Figure 11: Same results as for 10, but only including Admin 2 regions with more than 100 surveyed clusters. The Tweedie adjusted model is now practically unbiased (in regards to shrinkage), suggesting that the correction has larger issues with smaller samples.
Figure 12: Performance on a downstream causal inference task as a function of upstream model training epochs. Early training (e.g., after one epoch) yields poor model accuracy (high MAE) but minimal overfitting, allowing Tweedie’s correction to produce unbiased estimates using variance from either the calibration set ($\sigma_{\text{Cal.}}$) or training set ($\sigma_{\text{Train}}$). As epochs increase, model accuracy improves, but growing overfitting causes over-correction when using $\sigma_{\text{Cal.}}$, increasing bias, while $\sigma_{\text{Train}}$ maintains near-unbiased results.
@@
The DHS displaces survey coordinates ($\geq$ 3 km) to protect privacy, so we already address a built‑in corruption source, which shows improvement. Generally, a weak method or noisy data would make debiasing more critical, as uncertainty increases shrinkage. To show this, we ran an additional sensitivity analysis on the causal aid task (cf. Fig 6). We trained our model on the same dataset but shuffled a share of the labels in the train+cal to add noise, only increasing Tweedie’s relative advantage. Results are below in Tables 2 and 3.
@@
Table 2: Performance of the naive estimator under varying levels of label shuffling in training and calibration data.
@@
Table 3: Performance of Tweedie’s estimator under varying levels of label shuffling in training and calibration data.
Variance source for Tweedie. Figure 12 shows that estimating $\sigma$ on the training split (rather than the calibration split) mitigates over-correction at later epochs; our package therefore exposes both options, defaulting to $\sigma_{\text{Train}}$. We also present in 4 the results of an analysis probing the sensitivity of correction performance to cal set size (mean $\pm$ sd across 20 random subsamples).
@@
Table 4: Sensitivity of LCC and Tweedie’s estimators to calibration set size (mean $\pm$ sd across 20 random subsamples).
@@
(i) PPI (Angelopoulos et al. 2023) Fresh ground-truth labeled data at the downstream causal inference stage (new treatment-predictor-outcome tuples). None; treats model as black-box and applies rectifier term. Predictive model is fixed (not retrained); valid statistical coverage in large samples; independence assumptions for inference.
(ii) Ratledge et al. (2022) None additional; uses existing upstream training data. Yes; modifies loss function to include quintile-specific bias penalty ($\mathcal{L}=\mathrm{MSE}+\lambda_{r}\|\theta\|_{2}^{2}+\lambda_{b}\hat{E}_{b}$; if using AdamW, set $\lambda_{r}=0$). Requires retraining. Bias reducible by penalizing group-level discrepancies; assumes quintile-based error structure.
(iii) DSL (Egami et al. 2023) Gold-standard labeled data on a subset of units (no new data collection); known or estimated sampling probabilities depending on surrogates and covariates. None; post-hoc construction of bias-corrected pseudo-outcomes via cross-fitting and doubly-robust estimation. Known sampling design with probabilities bounded away from zero; double robustness (consistent if either outcome model or sampling model is correctly specified).
@@
(v) Tweedie’s correction (this work) Data from upstream for estimating $\sigma^{2}$ and score function (no new data collection). None; post-hoc adjustment using empirical Bayes and density scores. Gaussian homoskedastic noise ($\hat{Y}_{i}=Y_{i}-\varepsilon_{i}$, $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ independent of $Y_{i}$); $\mathbb{E}[\varepsilon_{i}\mid A_{i}]=0$.
Table 5: Characterization of Debiasing Methods for Causal Inference Using ML Predictions
Funder Sector Aid Sector Name $\widehat{\textrm{ATE}}^{Y}$ $\widehat{\textrm{ATE}}^{\hat{Y}}$ $\widehat{\textrm{ATE}}^{\hat{Y}}$, Tweedie # Treated # Control
@@
Table 6: Sector effect estimates, where the outcome scale is IWI units. Here, $\widehat{\textrm{ATE}}^{Y}$ refers to estimates of treatment effects using difference-in-means with DHS outcomes (considered to be ground truth). $\widehat{\textrm{ATE}}^{\hat{Y}}$ refers to estimates with predicted outcomes.
@@
Table 7: Overview of data: DHS surveys on health and living conditions.
Appendix D D. Simulation Design
@@
Appendix E E. World Bank and China Empirical Cases
For robustness, we repeat this experiment, using urban/rural, country, and survey year as confounders (Figure 13). The results show that this approach breaks the bias in terms of treatment effect. For larger studies, with more than 200 treated points, it also lowers the total error rate. We hypothesize that the reason the overall error rate does not decrease for small trials is due to the added variance introduced when calculating the Tweedie correction. The correction is a mean of the sample scores; it decreases with larger sample sizes.
Figure 13: Scatterplots comparing ATE estimated using predicted IWI versus observed IWI for aid interventions funded by China (red points) and the World Bank (blue points). Tweedie’s correction slightly improves the overall MAE.
Figure 14: Scatterplots comparing ATE estimates by funder when country, year and urban/rural status are used as confounders. Unlike in Figure 13, the samples were weighted using IPTW to account for these confounders.
Appendix F F. Implementing Strategies with unshrink
@@
For all-Africa IWI wealth map predictions with associated score and calibration parameters (e.g., $\sigma$) needed for adjustment:
@@
Figure 15: Results of the hyperparameter tuning for finding the best suited $\lambda_{b}$ for the Ratledge loss. Based of these results, we opted to use $\lambda_{b}=15$ as it had the least shrinkage bias $(b\approx 1)$.